OpenStudy (anonymous):

what is the answer of this equation? a^(2x)-2(a^x)=15

OpenStudy (anonymous):

\(\large a^{2x}-2a^x=15\) Let \(\large t=a^x\) we have: \(\large t^2-2t-15=0 \Rightarrow \begin{cases}t_1=5\\ t_2=-3 \end{cases}\) \(\large \color{red}{t_1=5} \Rightarrow a^x=5 \Rightarrow x=log_{a}{5}\) \(\large \color{red}{t_2=-3}\Rightarrow a^x=-3\Rightarrow x=log_{a}{-3}\)